Question 3 1 pts Find a particular solution of the equation y3/4=-4x2 2x - 8of the form yp(= Ax2 + Bx +C A = .C B =
Answer #1

y+3y+4y=-4x²+2x-8. let » → Up (x) = A x²+ Bx tc. Yp(x) = 2 AX + B Yp 072 : 2A. substituting above values in ① ☺ > 2A +6AX +3